OpenStudy (anonymous):

YZ has endpoints y(0,5) and z(12,3). Find the length of YZ to the nearest tenth.

OpenStudy (leader):

\(\ d(A,B) = \sqrt{(x_B - x_A)^2+(y_B - y_A)^2}\)

OpenStudy (anonymous):

would it be 12.2

OpenStudy (leader):

\[\begin{aligned} d(A,B) & = \sqrt{\left( 12 - 0 \right)^2 + \left( 3 - 5 \right)^2} \\ d(A,B) & = \sqrt{ 144 + 4} \\ d(A,B) & = 2 \sqrt{ 37 } \end{aligned}\]

OpenStudy (leader):

Yes it would be 12.2 if you round

OpenStudy (leader):

\(d(A,B) = 2 \sqrt{ 37 }~ \approx ~ 12.1655\)

OpenStudy (leader):

@jimmiedeel